OpenStudy (hdrager):

Water slowly evaporates from a circular shaped puddle. The radius of the puddle decreases at a rate of 7 in/hr. Assuming the puddle retains its circular shape, at what rate is the area of the puddle changing when the radius is 3 in?

OpenStudy (tkhunny):

You will need a formula for the area of a circle. You will need a derivative, in differential form. You are almost done.

OpenStudy (hdrager):

\[A=\pi r^2\]

OpenStudy (hdrager):

the derivative of the area formula?

OpenStudy (mrs.ambrose614):

A is 2

OpenStudy (hdrager):

wait why?

OpenStudy (hdrager):

@Mrs.ambrose614

OpenStudy (tkhunny):

\(A = \pi r^{2}\) Differential Form of the Derivative \(dA = 2\pi r\;dr\) Fill in the blanks.

OpenStudy (mrs.ambrose614):

A =2

OpenStudy (hdrager):

\[dA=-42\pi\]

OpenStudy (hdrager):

is that the answer?

OpenStudy (tkhunny):

No, that can't be the answer. Find your units and show your work. Good call. It SHOULD be negative, since it's shrinking.

OpenStudy (hdrager):

\[dA=2\pi r *dr\] \[dA=2\pi(3)*(-7)\] \[dA=-42\pi\]

OpenStudy (hdrager):

This is what I did

OpenStudy (tkhunny):

Well, that looks fine from a purely numerical point of view. Still needs units. Units will save you.

OpenStudy (hdrager):

oh so \[42\pi m^2/hr\]

OpenStudy (hdrager):

in not m sorry

OpenStudy (tkhunny):

\(2\cdot\pi\cdot (3\;in)\cdot \dfrac{7\;in}{hr} = -42\pi\;in^2/hr\)